OpenStudy (dimples97):

Simplify each expression. Use positive exponents. m3n–6p0 Help me solve please?

Nnesha (nnesha):

\[\large\rm m^3n^{-6} p^0\] like this ?

OpenStudy (dimples97):

yes

Nnesha (nnesha):

\[\huge\rm x^{:(} = \frac{1}{x^{:)}}\] flip the fraction to change negative to positive exponent

zepdrix (zepdrix):

lolol ^ :D

Nnesha (nnesha):

if negative was at the denominator \[\huge\rm \frac{ 1 }{ x^{:(} } =\frac{x^{:)}}{1} \large{~which ~is~same ~as= x^{:)}}\] i would flip the fraction to make it positive exponents and x/1 is same as xx

Nnesha (nnesha):

haha

OpenStudy (dimples97):

Hmmm... okay i see thanks!

Nnesha (nnesha):

so what would be the answer ?? hmm here is an example \[\huge\rm 2^{-4} = \frac{1}{2^4}\]

Nnesha (nnesha):

one more rule anything to the zero power is equal to one \[\rm (anything)^0 =1~~~~(Nnesha)^0 = 1 ~~(chocolates)^0=1\]